OpenStudy (anonymous):

How many solutions does a triangle with values a = 42, A = 117, and b = 34 have? Write your answer in numeric form.

OpenStudy (anonymous):

@Hero can you help?

hero (hero):

First find the measure of angle B using law of sines.

hero (hero):

\[\frac{\sin A}{a} = \frac{\sin B}{b}\]

hero (hero):

@julesstrawberry are you still here?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

is it 1 as a final answer? I called one of my friends and we worked through it, I just want to make sure I'm correct.

hero (hero):

Did you find the measure of angle B?

OpenStudy (anonymous):

yes

hero (hero):

Would you mind sharing?

OpenStudy (anonymous):

46.14

OpenStudy (anonymous):

would you mind helping out with my new question?

hero (hero):

Close this question first. Then post your next one.